### v1

# Gradient flow in parameter space is equivalent to linear interpolation in output space

Abstract.
We prove that the usual gradient flow in parameter space that underlies many training algorithms for neural networks in deep learning can be continuously deformed into an adapted gradient flow which yields (constrained) Euclidean gradient flow in output space. Moreover, if the Jacobian of the outputs with respect to the parameters is full rank (for fixed training data), then the time variable can be reparametrized so that the resulting flow is simply linear interpolation, and a global minimum can be achieved.
1. Introduction
At the core of most algorithms currently used for training neural networks is gradient descent, whose theoretical counterpart in the continuum limit is gradient flow. This flow is defined in a space of parameters $\mathbb{R}_{\underline{\theta}}^{K}$ with respect to a cost function defined in $\mathbb{R}_{\underline{x}}^{QN}$. The cost is non-convex as a function of the parameters, and standard gradient flow in $\mathbb{R}_{\underline{\theta}}^{K}$ might not converge to a global minimum. Even changing perspective to $\mathbb{R}_{\underline{x}}^{QN}$, where the cost can be chosen to be convex, the resulting associated flow is complicated.
In [Che23], it was discussed in detail how the choice of the Riemannian structure in $\mathbb{R}_{\underline{\theta}}^{K}$ influences the training dynamics. In this context, a modified gradient flow that induces the Euclidean gradient flow in output space $\mathbb{R}_{\underline{x}}^{QN}$ was introduced and contrasted with the standard Euclidean gradient flow in parameter space; it was shown that both flows exhibit the same critical sets. In the work at hand, we extend these results by proving that those two flows are in fact homotopy equivalent to one another (Theorem 2.3). Moreover, we prove that if the Jacobian matrix exhibits no rank loss, then the Euclidean flow in output space is reparametrization equivalent to linear interpolation with respect to a suitable time variable (Proposition 2.4). When there is rank loss, an expression for the deviation from linear interpolation is given (Proposition 2.5).
In section 3, we provide some applications: First, we note that instead of modifying the metric in $\mathbb{R}_{\underline{\theta}}^{K}$, one could prescribe a convenient path in $\mathbb{R}_{\underline{x}}^{QN}$ and then find an associated path in parameter space. Next, we show that neural collapse occurs in the output layer for the trivialized dynamics studied in 2.2 (Corollary 3.2). Finally, we rephrase Theorem 2.3 in terms of the neural tangent kernel.
We briefly comment on related work in section 4. A short appendix on generalized inverses for rank deficient matrices is provided as well.
2. Main results
Consider a family of functions
$f_{\underline{\theta}}:\mathbb{R}^{M}\to\mathbb{R}^{Q}$ (2.1)
parametrized by a vector of parameters $\underline{\theta}\in\mathbb{R}^{K}$. Suppose we wish to optimize $f_{\underline{\theta}}$ with respect to a given cost function $\mathcal{C}:\mathbb{R}^{K}\to\mathbb{R}$. A natural approach is to pick a starting point $\underline{\theta}_{0}\in\mathbb{R}^{K}$ and construct a path
$\underline{\theta}(s):[0,\infty)\to\mathbb{R}^{K}$ (2.2)
via the gradient flow
$\partial_{s}\underline{\theta}(s)$ $=-\nabla_{\underline{\theta}}\mathcal{C},$
$\underline{\theta}(0)$ $=\underline{\theta}_{0}.$ (2.3)
We will refer to (2) as the standard gradient flow.
When $f_{\underline{\theta}}$ is a neural network, the optimization problem has the following structure: Consider a set $\mathcal{X}_{0}\subset\mathbb{R}^{M}$ (the training data), such that $\left\lvert{\mathcal{X}_{0}}\right\rvert=N$. We can form a data matrix $X_{0}\in\mathbb{R}^{M\times N}$ by making the vectors in $\mathcal{X}_{0}$ into the columns of $X_{0}$. Then we can consider a function
$f(\underline{\theta},X_{0}):\mathbb{R}^{K}\times\mathbb{R}^{M\times N}\to\mathbb{R}^{Q\times N},$ (2.4)
defined by its action on each column of $X_{0}$: for $i=1,\cdots,N$,
$f(\underline{\theta},X_{0})_{i}=f_{\underline{\theta}}((X_{0})_{i})\in\mathbb{R}^{Q}.$ (2.5)
In the context of supervised learning, each data point $x\in\mathcal{X}_{0}$ is associated to a desired output $y(x)\in\mathbb{R}^{Q}$, and so from a data matrix $X_{0}$ we can form the corresponding matrix of labels, $Y\in\mathbb{R}^{Q\times N}$,
$Y_{i}=y((X_{0})_{i}).$ (2.6)
A common choice for cost function is the squared loss,
$\mathcal{C}(\underline{\theta},X_{0},Y)=\frac{1}{N}\left\lVert{f(\underline{\theta},X_{0})-Y}\right\rVert_{\mathcal{L}^{2}}^{2},$ (2.7)
where $\left\lVert{\cdot}\right\rVert_{\mathcal{L}^{2}}$ is the Frobenius or Hilbert-Schmidt norm.
For fixed training data $X_{0}=\left[x_{1}\cdots x_{j}\cdots x_{N}\right]$, we can also consider the function
$\underline{x}$ $:\mathbb{R}^{K}\to\mathbb{R}^{QN}$ (2.8)
from parameter space to output space defined by
$\underline{x}(\underline{\theta})$ $:=(f_{\underline{\theta}}(x_{1})^{T},\cdots,f_{\underline{\theta}}(x_{N})^{T})^{T}.$ (2.9)
We define the correspondent vector of labels $\underline{y}\in\mathbb{R}^{QN}$,
$\underline{y}:=(y(x_{1})^{T},\cdots,y(x_{N})^{T})^{T},$ (2.10)
In this case, the cost (2.7) becomes
$\mathcal{C}(\underline{x}(\underline{\theta}))=\frac{1}{N}\left\lvert{\underline{x}(\underline{\theta})-\underline{y}}\right\rvert^{2}.$ (2.11)
The Jacobian $D[\underline{\theta}]\in\mathbb{R}^{QN\times K}$ is
$\left(D[\underline{\theta}]\right)_{jk}$ $:=\frac{\partial\underline{x}_{j}[\underline{\theta}]}{\partial\theta_{k}},$ (2.12)
and we assume from now on that the partial derivatives in (2.12) exist and are Lipschitz continuous for all $\underline{\theta}\in\mathbb{R}^{K}$. We will often abbreviate $D=D[\underline{\theta}]$.
For $\underline{\theta}(s)\in\mathbb{R}^{K}$, we can consider the associated path
$\underline{x}(s):[0,\infty)\to\mathbb{R}^{QN}$ (2.13)
given by
$\underline{x}(s)$ $:=\underline{x}(\theta(s))$ (2.14)
from (2.9). The associated Jacobian is now also time dependent, and we will write $D(s):=D[\underline{\theta}(s)]$, or simply $D$ for convenience. The chain rule gives
$\partial_{s}\underline{x}(s)=D(s)\,\partial_{s}\underline{\theta}(s)$ (2.15)
and
$\nabla_{\underline{\theta}}\mathcal{C}=D^{T}\nabla_{\underline{x}}\mathcal{C}.$ (2.16)
Remark 2.1.
It is clear that
$\left\{\underline{\theta}:\nabla_{\underline{x}(\underline{\theta})}\mathcal{C}=0\right\}\subset\left\{\underline{\theta}:\nabla_{\underline{\theta}}\mathcal{C}=0\right\},$ (2.17)
however, the two sets are not necessarily identical. It is known that if $\operatorname{rank}D[\underline{\theta}]=QN=\min\{K,QN\}$, then $\nabla_{\underline{x}}\mathcal{C}=0$ if, and only if, $\nabla_{\underline{\theta}}\mathcal{C}=0$. Moreover, if $\mathcal{C}$ is the squared loss, then $\nabla_{\underline{x}}\mathcal{C}=0$ corresponds to a global minimum with zero loss. See e.g. [Che23, CE23, KMTM24] for more details.
2.1. Adapted gradient flow
By a simple computation using (2.15) and (2.16), if $\underline{\theta}(s)$ satisfies the standard gradient flow (2), then
$\partial_{s}\underline{x}=-DD^{T}\nabla_{\underline{x}}\mathcal{C}.$ (2.18)
We will refer to (2.18) as standard gradient flow in output space. Suppose instead we wished to modify $\underline{\theta}(s)$ so that the associated path $\underline{x}(s)$ satisfies the Euclidean gradient flow. We recall here results in previous work by one of the authors, in the overparametrized case:
Lemma 2.2 ([Che23]).
Let $K\geq QN$, and let $\underline{x}(s)=\underline{x}(\underline{\theta}(s))$ be defined as in (2.9). When $D$ is full rank, setting
$\partial_{s}\underline{\theta}(s)=-(D^{T}D){}^{+}\nabla_{\underline{\theta}}\mathcal{C}$ (2.19)
yields
$\partial_{s}\underline{x}(s)=-\nabla_{\underline{x}}\mathcal{C}.$ (2.20)
If we allow for $\operatorname{rank}(D)\leq QN$, then letting
$\partial_{s}\underline{\theta}(s)=-D^{T}\psi,$ (2.21)
for $\psi$ satisfying
$DD^{T}\psi=\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C},$ (2.22)
results in
$\partial_{s}\underline{x}(s)=-\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}.$ (2.23)
We will now condense these two separate cases into the same. First, note that
$\tilde{\psi}:=D(D){}^{+}(DD^{T}){}^{+}(D^{T}){}^{+}\nabla_{\underline{\theta}}\mathcal{C}$ (2.24)
satisfies (2.22), as the following computation shows:
[…]
$=DD^{T}(DD^{T}){}^{+}(D^{T}){}^{+}D^{T}\nabla_{\underline{x}}\mathcal{C}$
[…]
$=DD^{T}(DD^{T}){}^{+}\nabla_{\underline{x}}\mathcal{C}$
$=\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}.$
Moreover, we can rewrite
$-(D^{T}D){}^{+}\nabla_{\underline{\theta}}\mathcal{C}$ $=-(D^{T}D){}^{+}(D^{T}D)(D^{T}D){}^{+}\nabla_{\underline{\theta}}\mathcal{C}$
$=-(D^{T}D)(D^{T}D){}^{+}(D^{T}D){}^{+}\nabla_{\underline{\theta}}\mathcal{C}$
$=-D^{T}\left[D(D){}^{+}(D^{T}){}^{+}(D){}^{+}(D^{T}){}^{+}\right]\nabla_{\underline{\theta}}\mathcal{C}$ (2.26)
$=-D^{T}\left[D(D){}^{+}(DD^{T}){}^{+}(D^{T}){}^{+}\right]\nabla_{\underline{\theta}}\mathcal{C},$
$=-D^{T}\tilde{\psi}.$
Therefore in general, for $K\geq QN$ and $\operatorname{rank}(D)\leq QN$,
$\partial_{s}\underline{\theta}=-(D^{T}D){}^{+}\nabla_{\underline{\theta}}\mathcal{C}$ (2.27)
implies
$\partial_{s}\underline{x}=-\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}.$ (2.28)
We will refer to (2.27) as adapted gradient flow in $\mathbb{R}_{\underline{\theta}}^{K}$ and (2.28) as constrained Euclidean gradient flow in $\mathbb{R}_{\underline{x}}^{QN}$.
It was further shown in [Che23, Theorem 4.2] that the standard gradient flow (2) and the adapted gradient flow (2.27) have the same equilibrium points. We extend this to a family of interpolating gradient fields.
Theorem 2.3.
Let $\mathcal{C}:\mathbb{R}^{K}\to\mathbb{R}$ depend on $\underline{\theta}\in\mathbb{R}^{K}$ only through $\underline{x}(\underline{\theta})$, for $\underline{x}:\mathbb{R}^{K}\to\mathbb{R}^{QN}$ defined as in (2.9), and let $D[\underline{\theta}]\in\mathbb{R}^{QN\times K}$ be the Jacobian defined in (2.12). There is a one-parameter family of vector fields in $\mathbb{R}^{K}$ interpolating between
$-\nabla_{\underline{\theta}}\mathcal{C}$ (2.29)
and
$-(D^{T}D){}^{+}\nabla_{\underline{\theta}}\mathcal{C};$ (2.30)
this interpolation preserves singularities.
Moreover, this induces an interpolation between the vector fields for the standard gradient flow in output space (2.18) and the constrained Euclidean gradient flow (2.28) which preserves the equilibrium points of these flows.
Proof.
Define the family of vector fields
$V_{\underline{\theta},\alpha}:=-A_{\underline{\theta},\alpha}\nabla_{\underline{\theta}}\mathcal{C},\quad\alpha\in[0,1],$ (2.31)
for
$A_{\underline{\theta},\alpha}:=\left(\alpha(D^{T}D){}^{+}+(1-\alpha){\textbf{1}}_{K\times K}\right).$ (2.32)
Then
$V_{\underline{\theta},0}=-\nabla_{\underline{\theta}}\mathcal{C}$ (2.33)
and
$V_{\underline{\theta},1}=-(D^{T}D){}^{+}\nabla_{\underline{\theta}}\mathcal{C}.$ (2.34)
For $\alpha\in[0,1)$, note that $A_{\underline{\theta},\alpha}>0$ for all $\underline{\theta}$. For $\alpha=1$, writing $\mathcal{P}_{\operatorname{ran}(D^{T})}=(D){}^{+}D$, we have
$\nabla_{\underline{\theta}}\mathcal{C}=(D){}^{+}D\nabla_{\underline{\theta}}\mathcal{C},$ (2.35)
since (2.16) shows $\nabla_{\underline{\theta}}\mathcal{C}\in\operatorname{ran}(D^{T})$, and also
$\ker(D^{T}D){}^{+}=\ker(D^{T}D)^{T}=\ker(D^{T}D)=\ker((D){}^{+}D),$ (2.36)
where we used (A.11) and Lemma A.2. Thus, $V_{\underline{\theta},\alpha}=0$ if, and only if, $\nabla_{\underline{\theta}}\mathcal{C}=0$, and the vector fields $V_{\underline{\theta},\alpha}$ are singular at the same points $\underline{\theta}\in\mathbb{R}^{K}$, for all $\alpha\in[0,1]$.
Suppose now that
$\partial_{s}\underline{\theta}(s)=V_{\underline{\theta}(s),\alpha},$ (2.37)
so that for $\underline{x}(s)=\underline{x}(\underline{\theta}(s))$,
$\partial_{s}\underline{x}(s)=D[\underline{\theta}(s)]V_{\underline{\theta}(s),\alpha}.$ (2.38)
We wish to write $\partial_{s}\underline{x}$ in terms of $\nabla_{\underline{x}}\mathcal{C}$, so we compute
$D\,V_{\underline{\theta},\alpha}$ $=-D\left(\alpha(D^{T}D){}^{+}+(1-\alpha){\textbf{1}}_{K\times K}\right)\nabla_{\underline{\theta}}\mathcal{C}$
$=-\left(\alpha D(D^{T}D){}^{+}+(1-\alpha)D\right)D^{T}\nabla_{\underline{x}}\mathcal{C}$
$=-\left(\alpha D(D^{T}D){}^{+}D^{T}+(1-\alpha)DD^{T}\right)\nabla_{\underline{x}}\mathcal{C}$ (2.39)
=$\stackrel{{\scriptstyle(*)}}{{=}}-\left(\alpha\mathcal{P}_{\operatorname{ran}(DD^{T})}+(1-\alpha)DD^{T}\right)\nabla_{\underline{x}}\mathcal{C}$
$=-\left(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)DD^{T}\right)\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C},$
where ()(*)( ∗ ) follows from (A.7), (A.8) and (A.12). Therefore,
$\partial_{s}\underline{x}(s)=V_{\underline{x}(s),\alpha}$ (2.40)
for vector fields in $\mathbb{R}^{QN}$ defined along the path $\underline{x}(s)$,
$V_{\underline{x}(s),\alpha}:=-\left(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)D(s)D(s)^{T}\right)\mathcal{P}_{\operatorname{ran}(D(s)D(s)^{T})}\nabla_{\underline{x}}\mathcal{C}.$ (2.41)
Observe that
$V_{\underline{\theta}(s),0}=-\nabla_{\underline{\theta}}\mathcal{C}$ (2.42)
and
$V_{\underline{x}(s),1}=-\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C},$ (2.43)
so that in fact $V_{\underline{x}(s),\alpha},\,\alpha\in[0,1],$ interpolates between the standard dynamics given by (2) at $\alpha=0$, and constrained Euclidean gradient flow (2.28) for $\underline{x}(s)$ at $\alpha=1$. Moreover, for $\alpha\in(0,1]$, $(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)DD^{T})$ is positive-definite, and for $\alpha=0$, $DD^{T}$ is injective in the range of $\mathcal{P}_{\operatorname{ran}(DD^{T})}$, so
$(\alpha{\textbf{1}}+(1-\alpha)DD^{T})\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}=0\text{ if, and only if, }\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}=0.$ (2.44)
Thus, the equilibrium points of (2.40) are the same for all $\alpha\in[0,1]$. ∎
We add the following remarks highlighting the geometric interpretation of the above results. If $DD^{T}$ has full rank $QN$, then $V_{\underline{x}(s),\alpha}$ is the gradient vector field with respect to the Riemannian metric on $T\mathbb{R}^{QN}$ with tensor
$\left(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)DD^{T}\right)^{-1}\,.$ (2.45)
This means that $V_{\underline{x}(s),\alpha}$ can equivalently be considered as the gradient fields obtained from the family of Riemannian structures interpolating between the Euclidean structure on $T\mathbb{R}^{QN}$, and the metric structure induced by the Euclidean structure on parameter space.
If $DD^{T}$ is rank-deficient, $\operatorname{rank}(DD^{T})<QN$, we let $\mathcal{V}\subset T\mathbb{R}^{QN}$ denote the vector subbundle whose fibers are given by the range of $DD^{T}$. In this situation, $V_{\underline{x}(s),\alpha}$ is the gradient with respect to the bundle metric $h_{\alpha}$ on $\mathcal{V}$ with tensor
$[h_{\alpha}]=\left(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)DD^{T}\right)^{-1}\Big{|}_{\operatorname{ran}(DD^{T}[\underline{\theta}])}\,.$ (2.46)
The triple $(\mathbb{R}^{QN},\mathcal{V},h_{\alpha})$ defines a sub-Riemannian manifold with a family of bundle metrics for $\alpha\in[0,1]$; see [Che23] for a related discussion and further details.
2.2. Reparametrization
Assuming $\operatorname{rank}{D}=QN\leq K$, the adapted gradient flow in output space resulting from (2.27) is simply Euclidean gradient flow,
$\partial_{s}\underline{x}(s)$ $=-\nabla_{\underline{x}}\mathcal{C}(\underline{x}(s)).$ (2.47)
For
$\mathcal{C}(\underline{x}(\underline{\theta}))=\frac{1}{2N}\left\lvert{\underline{x}(\underline{\theta})-\underline{y}}\right\rvert_{\mathbb{R}^{QN}}^{2}$ (2.48)
we have
$\partial_{s}\underline{x}=\frac{1}{N}(\underline{x}(s)-\underline{y}),$ (2.49)
which is solved by
$\underline{x}(s)=\underline{y}+e^{-\frac{s}{N}}(\underline{x}(0)-\underline{y}),$ (2.50)
and we see that $\underline{x}(s)\to\underline{y}$ as $s\to\infty$.
Suppose instead that we wanted to impose a linear interpolation
$\underline{\hat{x}}(t)$ $=\underline{y}+(1-t)(\underline{x}_{0}-\underline{y}).$ (2.51)
Then letting
$t$ $:=1-e^{-\frac{s}{N}},$ (2.52)
$s$ $=-N\ln(1-t),$ (2.53)
and
$\underline{\tilde{x}}(t):=\underline{x}(-N\ln(1-t)),$ (2.54)
for $\underline{x}(s)$ a solution to (2.47), we have
$\partial_{t}\underline{\tilde{x}}(t)$ $=\partial_{t}\underline{x}(-N\ln(1-t))$
$=-\nabla_{\underline{x}}\mathcal{C}[\underline{\tilde{x}}(t)]\frac{N}{1-t}$
$=-\frac{1}{1-t}(\underline{\tilde{x}}(t)-\underline{y}).$ (2.55)
Note that $\underline{\hat{x}}(t)$ solves (2.2), as
$\partial_{t}\underline{\hat{x}}(t)$ $=-(\underline{\hat{x}}_{0}-\underline{y})$
$=-\frac{1}{1-t}(\underline{\hat{x}}(t)-\underline{y}).$ (2.56)
Thus, we have proved the following:
Proposition 2.4.
If $\underline{x}(s)$ is a solution to Euclidean gradient flow (2.47), then
$\underline{\tilde{x}}(t):=\underline{x}(-N\ln(1-t))=\underline{y}+(1-t)(\underline{x}_{0}-\underline{y}),$ (2.57)
and $\underline{\tilde{x}}(t)\to\underline{y}$ as $t\to 1$. In particular, this holds for $\underline{x}(s)=\underline{x}(\underline{\theta}(s))$ defined as in (2.9) when $\underline{\theta}(s)$ is a solution to the adapted gradient flow (2.27) and $\operatorname{rank}D[\underline{\theta}(s)]=QN\leq K$.
On the other hand, if $\operatorname{rank}(D)<QN$, the reparametrized Euclidean gradient flow in output space provides a concrete criterion whereby the effect of rank loss can be measured.
Proposition 2.5.
In general, the constrained Euclidean gradient flow in output space with time reparametrization as in (2.54) satisfies
$\partial_{t}\underline{\tilde{x}}(t)=-\frac{1}{1-t}\mathcal{P}_{t}(\underline{\tilde{x}}(t)-\underline{y})\,,\;\;\;\underline{\tilde{x}}(0)=\underline{x}_{0}\,,\;\;\;t\in[0,1)\,,$ (2.58)
where $\underline{\tilde{x}}(t)=\underline{x}[\underline{\theta}(s(t))]$ and $\mathcal{P}_{t}:=\mathcal{P}_{\operatorname{ran}(DD^{T})}[\underline{\theta}(s(t))]$. When $\operatorname{rank}(D)<QN$, the deviation from linear interpolation is given by
[…]
where the linear propagator $\mathcal{U}_{t,t^{\prime}}$ is determined by
$\partial_{t}\mathcal{U}_{t,t^{\prime}}=\frac{1}{1-t}\,\mathcal{P}_{t}\,\mathcal{U}_{t,t^{\prime}}\,,\;\;\;\mathcal{U}_{t^{\prime},t^{\prime}}={\textbf{1}}_{QN\times QN}\,,\;\;\;t,t^{\prime}\in[0,1)\,.$ (2.60)
Proof.
The equation (2.58) is straightforwardly obtained in a similar way as (2.2).
To prove (2.59), we write
$\underline{\tilde{x}}(t)=(1-t)\underline{x}_{0}+t\underline{y}+(1-t)R(t)$ (2.61)
so that
$\partial_{t}\underline{\tilde{x}}(t)=-(\underline{x}_{0}-\underline{y})-R(t)+(1-t)\partial_{t}R(t)\,.$ (2.62)
On the other hand,
$\partial_{t}\underline{\tilde{x}}(t)$ $=-\frac{1}{1-t}\mathcal{P}_{t}(\underline{\tilde{x}}(t)-\underline{y})$
$=-\frac{1}{1-t}(\underline{\tilde{x}}(t)-\underline{y})+\frac{1}{1-t}\mathcal{P}_{t}^{\perp}(\underline{\tilde{x}}(t)-\underline{y})$ (2.63)
$=-\frac{1}{1-t}\big{(}(1-t)\underline{x}_{0}+t\underline{y}-\underline{y}+(1-t)R(t)\big{)}+\frac{1}{1-t}\mathcal{P}_{t}^{\perp}\big{(}(1-t)\underline{x}_{0}+t\underline{y}-\underline{y}+(1-t)R(t))\big{)}$
$=-(\underline{x}_{0}-\underline{y})-R(t)+\mathcal{P}_{t}^{\perp}\big{(}\underline{x}_{0}-\underline{y}+R(t)\big{)},$
where we used (2.61) to pass to the third line. Comparing the last line with (2.62), we find
$\partial_{t}R(t)=\frac{1}{1-t}\mathcal{P}_{t}^{\perp}R(t)+\frac{1}{1-t}\mathcal{P}_{t}^{\perp}(\underline{x}_{0}-\underline{y})\,.$ (2.64)
Solving this matrix valued linear ODE for $R(t)$ using the Duhamel (or variation of constants) formula, and substituting the resulting expression in (2.61) yields (2.59), as claimed. ∎
3. Applications
3.1. Prescribed paths in output space
In the previous section, we discussed how changes in $\partial_{s}\underline{\theta}(s)$ influence $\partial_{s}\underline{x}(s)$. It could be interesting to take a different point of view and find $\underline{\theta}(s)$ from a given path $\underline{x}(s)$. In general, from (2.15),
$\mathcal{P}_{\operatorname{ran}(D^{T})}\partial_{s}\underline{\theta}=(D){}^{+}\partial_{s}\underline{x},$ (3.1)
and so assuming $\partial_{s}\underline{\theta}=\mathcal{P}_{\operatorname{ran}(D^{T})}\partial_{s}\underline{\theta}$, [^1] Note that this is satisfied by the standard gradient flow $\partial_{s}\underline{\theta}=-\nabla_{\underline{\theta}}\mathcal{C}=-D^{T}\nabla_{\underline{x}}\mathcal{C}.$ we can write
$\partial_{s}\underline{\theta}(s)=(D){}^{+}\partial_{s}\underline{x}(s).$ (3.2)
Thus, we could prescribe a path
$\underline{x}(s),$ $\quad s\in[0,T],$
$\partial_{s}\underline{x}(s)$ $\in\operatorname{ran}(D(s)),$ (3.3)
$\underline{x}(T)$ $=\underline{y},$
and then search for a corresponding $\underline{\theta}_{*}$ which realizes
$\underline{x}(\underline{\theta}_{*})=\underline{x}(T)$ (3.4)
by solving
$\partial_{s}\underline{\theta}(s)$ $=(D(s)){}^{+}\partial_{s}\underline{x}(s),$
$\underline{\theta}(0)$ $=\underline{\theta}_{0},$ (3.5)
[…]
for some initialization $\underline{\theta}_{0}\in\mathbb{R}^{K}$. Observe that
$\operatorname{ran}(D)=(\ker(D^{T}))^{\perp}=(\ker(D){}^{+})^{\perp},$ (3.6)
and so all equilibrium points of (3.1) satisfy $\partial_{s}\underline{x}=0$.
As an example, motivated by the results in section 2.2, one could take $\underline{x}(s)$ to be the linear interpolation (2.51) and check if, for a given $\underline{\theta}_{0}$,
$(\underline{x}_{0}-\underline{y})\in\operatorname{ran}{D(s)},$ (3.7)
for all $s\in[0,T]$.
Remark 3.1.
Whether or not a path $\underline{x}(s)$ satisfies the constraint in (3.1) (or (3.7) in the example above) could depend on the choice of $\underline{\theta}_{0}$. In general, properties of the Jacobian matrix $D[\underline{\theta}(s)]$ that determine the dynamics of $\underline{x}(s)$, such as rank and the subspace $\operatorname{ran}(DD^{T})$, might depend on the initialization even within a class $\left\{\underline{\theta}_{0}:\underline{x}(\underline{\theta}_{0})=\underline{x}_{0}\right\}$.
3.2. Final layer collapse
We turn to the case of a classification task. Assume without any loss of generality that the components of $\underline{\tilde{x}}$ are ordered by way of
$\underline{\tilde{x}}=(\dots\dots,\tilde{x}_{j,1}^{T},\dots,\tilde{x}_{j,i_{j}}^{T},\dots,\tilde{x}_{j,N_{j}}^{T},\dots\dots)^{T}\in\mathbb{R}^{QN}$ (3.8)
where each $\tilde{x}_{j,i_{j}}\in\mathbb{R}^{Q}$ belongs to the class labeled by $y_{j}\in\mathbb{R}^{Q}$, for $j=1,\dots,Q$, $i_{j}=1,\dots,N_{j}$, and $\sum_{j=1}^{Q}N_{j}=N$. Then, defining the class averages
$\overline{\tilde{x}_{j}}(t):=\frac{1}{N_{j}}\sum_{i_{j}=1}^{N_{j}}\tilde{x}_{j,i_{j}}(t)$ (3.9)
and the deviations
$\Delta\tilde{x}_{j,i_{j}}(t):=\tilde{x}_{j,i_{j}}(t)-\overline{\tilde{x}_{j}}(t)\,,$ (3.10)
the $\mathcal{L}^{2}$ cost (2.7) decomposes into
$\mathcal{C}[\underline{\tilde{x}}(t)]$ $=\frac{1}{N}\sum_{j=1}^{Q}\sum_{i_{j}}^{N_{j}}|\tilde{x}_{j,i_{j}}(t)-y_{j}|^{2}$
$=\frac{1}{N}\sum_{j=1}^{Q}\sum_{i_{j}}^{N_{j}}\big{(}|\tilde{x}_{j,i_{j}}(t)-\overline{\tilde{x}_{j}}(t)|^{2}+|\overline{\tilde{x}_{j}}(t)-y_{j}|^{2}\big{)}$
$=\frac{1}{N}\sum_{j=1}^{Q}\sum_{i_{j}}^{N_{j}}|\Delta\tilde{x}_{j,i_{j}}(t)|^{2}+\sum_{j=1}^{Q}|\overline{\tilde{x}_{j}}(t)-y_{j}|^{2}\,.$ (3.11)
This implies that zero loss training is achievable if and only if the class averages in the output layer are matched to the reference outputs, $\overline{\tilde{x}_{j}}(t)\rightarrow y_{j}$ for all $j=1,\dots,Q$, and all deviations vanish, $\Delta\tilde{x}_{j,i_{j}}(t)\rightarrow 0$ for all $j=1,\dots,Q$, $i_{j}=1,\dots,N_{j}$. The latter implies that the image of training data in the output layer contracts to one point per class.
Thus, Proposition 2.4 allows us to arrive at the following conclusion.
Corollary 3.2.
If $\operatorname{rank}(D)=QN$, then zero loss minimization is achieved, whereby the cluster averages converge to the reference outputs,
$\lim_{t\rightarrow 1^{-}}\overline{\tilde{x}_{j}}(t)=y_{j},$ (3.12)
and the deviations converge to zero,
$\lim_{t\rightarrow 1^{-}}\Delta{\tilde{x}}_{j,i_{j}}(t)=0\,,$ (3.13)
for all $j=1,\dots,Q$ and $i_{j}=1,\dots,N_{j}$.
The decomposition of the cost function (3.2) is related to the phenomenon known as neural collapse ([PHD20, HPD22]), which takes place on the penultimate layer of a neural network. For work on the relationship between neural collapse and final layer collapse, see e.g. [EW22].
3.3. Tangent kernel
A related point of view is to study the neural tangent kernel (NTK) introduced by [JGH18]. It can be written as
$\Theta(x,y,\underline{\theta})$ $=Df_{\underline{\theta}}(x)(Df_{\underline{\theta}}(y))^{T}\in\mathbb{R}^{Q\times Q},$ (3.14)
for a function $f_{\underline{\theta}}:\mathbb{R}^{M}\to\mathbb{R}^{Q}$ and $Df_{\underline{\theta}}(x)$ the Jacobian of $f_{\underline{\theta}}$ at $x\in\mathbb{R}^{M}$. Then for a set of training data $\mathcal{X}_{0}=\{x_{1},\cdots,x_{N}\}$ and $D[\underline{\theta}]$ defined as in section 2 we can write $DD^{T}\in\mathbb{R}^{QN\times QN}$ as a block matrix with blocks of size $Q\times Q$,
$(D[\underline{\theta}]D[\underline{\theta}]^{T})i,j$ $=\Theta(x_{i},x_{j},\underline{\theta}),\quad i,j=1,\cdots,N.$ (3.15)
The kernel $\Theta(x,y,\underline{\theta})$ is positive-definite with respect to the chosen training data at $\underline{\theta}$ if,  [^2] This is stronger than the definition given in [JGH18], which gives a similar statement for all functions in a given function class $\mathcal{F}$. for all $(u_{i})_{i=1}^{N}\in\bigoplus^{N}\mathbb{R}^{Q}$,
$\sum_{i=1}^{N}\left\lvert{u_{i}}\right\rvert^{2}>0\text{ implies }\sum_{i,j=1}^{N}u_{i}^{T}\Theta(x_{i},x_{j},\underline{\theta})u_{j}>0\,.$ (3.16)
From (3.15) we see that his holds if, and only if, $DD^{T}$ is a positive-definite matrix.
Thus, Theorem 2.3 suggests that, by changing the metric in parameter space, the NTK can be modified into a simpler kernel $K$ which preserves the equilibrium points of the flow. When $\Theta$ is positive-definite, $K$ satisfies
$K(x_{i},x_{j})=\delta_{ij}{\textbf{1}}_{Q\times Q},$ (3.17)
so that it is constant throughout training at least when evaluated on $\mathcal{X}_{0}$. This trivializes the training dynamics in output space, and holds for any neural network, or indeed any family of functions parametrized by $\underline{\theta}$.
4. Some related work
Linear paths
Some works, starting with [GVS15], interpolate the initial parameters $\underline{\theta}_{0}$ and final parameters $\underline{\theta}_{f}$ obtained after training the network with usual methods such as stochastic gradient descent,
$\underline{\theta}_{\alpha}=(1-\alpha)\underline{\theta}_{0}+\alpha\underline{\theta}_{f},\,\alpha\in[0,1],$ (4.1)
and study the behavior of the loss function on these linear slices of parameter space. Recent works, such as [LBZ+21], provide empirical and theoretical investigations of the monotonic linear interpolation (MLI) property, whereby the loss decreases monotonically along the path $\underline{\theta}_{\alpha}$. This is a different approach from the one taken in this paper, where the linear interpolation happens in output space.
In [GKRBZ24], the authors find linear trajectories of probability distributions when studying generative models and modifying the cost function.
The Jacobian (2.12)
As already discussed in subsection 3.3, the neural tangent kernel introduced by [JGH18] is closely related to the Jacobian studied in this paper. It inspired works on linear approximations of gradient flow/descent and the so called kernel or lazy regime, see e.g. [COB19, WGL+20].
Certain works study the rank of $D$ more explicitly: In [GLMW22], the authors study the functional dimension of ReLU neural networks, defined in terms of the rank of the Jacobian. In [KMTM24], the authors show that for mildly overparametrized ReLU neural networks with scalar output ($Q=1$) most activation patterns correspond to regions in parameter space where the Jacobian is full rank.
Acknowledgments: T.C. gratefully acknowledges support by the NSF through the grant DMS-2009800, and the RTG Grant DMS-1840314 - Analysis of PDE. P.M.E. was supported by NSF grant DMS-2009800 through T.C., and UT Austin’s GS Summer fellowship.
Appendix A Pseudoinverse
We recall here the definition and properties of the pseudoinverse of a (possibly rank-deficient) matrix. For proofs and more detail, we refer the reader to [CM09], chapter 1.
Definition A.1.
The pseudoinverse of $A\in\mathbb{R}^{m\times n}$ is the unique matrix $(A){}^{+}\in\mathbb{R}^{n\times m}$ such that
$A(A){}^{+}A$ $=A,$ (A.1)
$(A){}^{+}A(A){}^{+}$ $=(A){}^{+},$ (A.2)
$(A(A){}^{+})^{T}$ $=A(A){}^{+},$ (A.3)
$((A){}^{+}A)^{T}$ $=(A){}^{+}A.$ (A.4)
When $A$ is full rank, there are simple expressions for the pseudoinverse: When $\operatorname{rank}(A)=m=\min\{m,n\}$, $A$ is surjective, so $AA^{T}$ is invertible and we have $(A){}^{+}=A^{T}(AA^{T})^{-1}$. On the other hand, when $\operatorname{rank}(A)=n=\min\{m,n\}$, $A$ is injective and $A^{T}A$ is invertible and we have $(A){}^{+}=(A^{T}A)^{-1}A^{T}$. In general, there exist many algorithms to construct the pseudoinverse, for instance using singular value decomposition; see e.g. [CM09], chapter 12.
The matrices $A$ and $(A){}^{+}$ as above satisfy a few basic properties:
$((A){}^{+}){}^{+}$ $=A,$ (A.5)
$(A^{T}){}^{+}$ $=((A){}^{+})^{T},$ (A.6)
$(A^{T}A){}^{+}$ $=(A){}^{+}(A^{T}){}^{+}.$ (A.7)
Moreover, the pseudoinverse gives a convenient way to write projectors: Consider a subspace $V\subset\mathbb{R}^{d}$. We will denote by $\mathcal{P}_{V}$ the orthogonal projector onto $V$, which satisfies $\mathcal{P}_{V}=\mathcal{P}_{V}^{2}=\mathcal{P}_{V}^{T}$. Then
$A(A){}^{+}$ $=\mathcal{P}_{\operatorname{ran}(A)},$ (A.8)
$(A){}^{+}A$ $=\mathcal{P}_{\operatorname{ran}(A^{T})}.$ (A.9)
When considering the subspaces associated to $A$, it is useful to know the relations
$\operatorname{ran}(A){}^{+}$ $=\operatorname{ran}A^{T},$ (A.10)
$\ker(A){}^{+}$ $=\ker A^{T}.$ (A.11)
The following is a useful identity:
$A=AA^{T}(A^{T}){}^{+}=(A^{T}){}^{+}A^{T}A.$ (A.12)
Similar identities can be found by substituting $A^{T}$ and $(A){}^{+}$ in place of $A$.
Finally, note that it is not true in general that $(A){}^{+}A=A(A){}^{+}$; a sufficient condition is $A^{T}=A$. We make use of this fact, along with the following lemma, in the main text.
Lemma A.2.
$\ker(A^{T}A)=\ker((A){}^{+}A)$.
Proof.
We expand
$A^{T}A$ $=A^{T}A(A){}^{+}A,$ (A.13)
so
$\ker((A){}^{+}A)\subset\ker(A^{T}A).$ (A.14)
Similarly,
[…]
and so
$\ker(A^{T}A)\subset\ker((A){}^{+}A).$ (A.16)

### v2

@@
 Abstract.
-We prove that the usual gradient flow in parameter space that underlies many training algorithms for neural networks in deep learning can be continuously deformed into an adapted gradient flow which yields (constrained) Euclidean gradient flow in output space.
-Moreover, if the Jacobian of the outputs with respect to the parameters is full rank (for fixed training data), then the time variable can be reparametrized so that the resulting flow is simply linear interpolation, and a global minimum can be achieved.
+We prove that the standard gradient flow in parameter space that underlies many training algorithms in deep learning can be continuously deformed into an adapted gradient flow which yields (constrained) Euclidean gradient flow in output space.
+Moreover, for the $L^{2}$ loss, if the Jacobian of the outputs with respect to the parameters is full rank (for fixed training data), then the time variable can be reparametrized so that the resulting flow is simply linear interpolation, and a global minimum can be achieved.
+For the cross-entropy loss, under the same rank condition and assuming the labels have positive components, we derive an explicit formula for the unique global minimum.
 1. Introduction
@@
 In the work at hand, we extend these results by proving that those two flows are in fact homotopy equivalent to one another (Theorem 2.3).
-Moreover, we prove that if the Jacobian matrix exhibits no rank loss, then the Euclidean flow in output space is reparametrization equivalent to linear interpolation with respect to a suitable time variable (Proposition 2.4).
+Moreover, for the $L^{2}$ cost, we prove that if the Jacobian matrix exhibits no rank loss, then the Euclidean flow in output space is reparametrization equivalent to linear interpolation with respect to a suitable time variable (Proposition 2.4).
 When there is rank loss, an expression for the deviation from linear interpolation is given (Proposition 2.5).
+For the cross-entropy loss, we prove that output space fibers into invariant manifolds over which Euclidean gradient flow converges to a unique global minimum (Proposition 2.7).
 In section 3, we provide some applications: First, we note that instead of modifying the metric in $\mathbb{R}_{\underline{\theta}}^{K}$, one could prescribe a convenient path in $\mathbb{R}_{\underline{x}}^{QN}$ and then find an associated path in parameter space.
@@
 Finally, we rephrase Theorem 2.3 in terms of the neural tangent kernel.
-We briefly comment on related work in section 4. A short appendix on generalized inverses for rank deficient matrices is provided as well.
+We briefly comment on related work in section 4. A short appendix on pseudoinverses for rank deficient matrices is provided as well.
 2. Main results
@@
 $Y_{i}=y((X_{0})_{i}).$ (2.6)
-A common choice for cost function is the squared loss,
-$\mathcal{C}(\underline{\theta},X_{0},Y)=\frac{1}{N}\left\lVert{f(\underline{\theta},X_{0})-Y}\right\rVert_{\mathcal{L}^{2}}^{2},$ (2.7)
-where $\left\lVert{\cdot}\right\rVert_{\mathcal{L}^{2}}$ is the Frobenius or Hilbert-Schmidt norm.
+The total cost $\mathcal{C}$ used for training is usually a sum of errors for each data point,
+$\mathcal{C}(\underline{\theta})=\sum_{x\in\mathcal{X}_{0}}\ell(f(\underline{\theta},x),y(x)),$ (2.7)
+where the chosen loss function $\ell:\mathbb{R}^{Q}\times\mathbb{R}^{Q}\to\mathbb{R}$ depends on the parameters only through $f(\underline{\theta},x).$
 For fixed training data $X_{0}=\left[x_{1}\cdots x_{j}\cdots x_{N}\right]$, we can also consider the function
@@
 We define the correspondent vector of labels $\underline{y}\in\mathbb{R}^{QN}$,
-$\underline{y}:=(y(x_{1})^{T},\cdots,y(x_{N})^{T})^{T},$ (2.10)
-In this case, the cost (2.7) becomes
-$\mathcal{C}(\underline{x}(\underline{\theta}))=\frac{1}{N}\left\lvert{\underline{x}(\underline{\theta})-\underline{y}}\right\rvert^{2}.$ (2.11)
+$\underline{y}:=(y(x_{1})^{T},\cdots,y(x_{N})^{T})^{T}.$ (2.10)
+In this case the cost can be redefined as $\mathcal{C}(\underline{x}(\underline{\theta}))$, and only depends on the parameters through its dependence on $\underline{x}(\underline{\theta})$.
 The Jacobian $D[\underline{\theta}]\in\mathbb{R}^{QN\times K}$ is
@@
 given by
-$\underline{x}(s)$ $:=\underline{x}(\theta(s))$ (2.14)
+$\underline{x}(s)$ $:=\underline{x}(\underline{\theta}(s))$ (2.13)
 from (2.9). The associated Jacobian is now also time dependent, and we will write $D(s):=D[\underline{\theta}(s)]$, or simply $D$ for convenience. The chain rule gives
@@
 See e.g.
-[Che23, CE23, KMTM24] for more details.
+[Coo18, TKB19, Che23, KMTM24, CE25] for more details.
+Notation.
+Given a matrix $A\in\mathbb{R}^{m\times n}$, we denote by $(A){}^{+}$ its pseudoinverse, or Moore–Penrose inverse; see appendix A for a definition and useful properties.
+Moreover, given a subspace $V\subset\mathbb{R}^{d}$, we denote by $\mathcal{P}_{V}$ the orthogonal projection onto $V$, such that $\mathcal{P}_{V}=\mathcal{P}_{V}^{2}=\mathcal{P}_{V}^{T}$.
 2.1. Adapted gradient flow
@@
 Lemma 2.2 ([Che23]).
-Let $K\geq QN$, and let $\underline{x}(s)=\underline{x}(\underline{\theta}(s))$ be defined as in (2.9). When $D$ is full rank, setting
+Let $K\geq QN$, and let $\underline{x}(s)=\underline{x}(\underline{\theta}(s))$ be defined as in (2.9).
+Assume $\underline{x}(\underline{\theta})$ has Lipschitz continuous derivatives.
+When $D$ is full rank, setting
 $\partial_{s}\underline{\theta}(s)=-(D^{T}D){}^{+}\nabla_{\underline{\theta}}\mathcal{C}$ (2.18)
@@
 implies
-$\partial_{s}\underline{x}=-\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}.$ (2.28)
+$\partial_{s}\underline{x}=-\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}=-\mathcal{P}_{\operatorname{ran}(D)}\nabla_{\underline{x}}\mathcal{C}.$ (2.27)
 We will refer to (2.26) as adapted gradient flow in $\mathbb{R}_{\underline{\theta}}^{K}$ and (2.27) as constrained Euclidean gradient flow in $\mathbb{R}_{\underline{x}}^{QN}$.
@@
 Theorem 2.3.
-Let $\mathcal{C}:\mathbb{R}^{K}\to\mathbb{R}$ depend on $\underline{\theta}\in\mathbb{R}^{K}$ only through $\underline{x}(\underline{\theta})$, for $\underline{x}:\mathbb{R}^{K}\to\mathbb{R}^{QN}$ defined as in (2.9), and let $D[\underline{\theta}]\in\mathbb{R}^{QN\times K}$ be the Jacobian defined in (2.12).
+Let $\mathcal{C}:\mathbb{R}^{K}\to\mathbb{R}$ depend on $\underline{\theta}\in\mathbb{R}^{K}$ only through $\underline{x}(\underline{\theta})$, for $\underline{x}:\mathbb{R}^{K}\to\mathbb{R}^{QN}$ a function with Lipschitz continuous derivatives defined as in (2.9), and let $D[\underline{\theta}]\in\mathbb{R}^{QN\times K}$ be the Jacobian defined in (2.11).
 There is a one-parameter family of vector fields in $\mathbb{R}^{K}$ interpolating between
@@
 Moreover, this induces an interpolation between the vector fields for the standard gradient flow in output space (2.17) and the constrained Euclidean gradient flow (2.27) which preserves the equilibrium points of these flows.
+This holds for any loss function $\ell$ in (2.7), and to any neural network whose activation function has Lipschitz continuous derivatives.
+Non-smooth activation functions, such as ReLU, can be mollified to satisfy this condition.
 Proof.
@@
 The triple $(\mathbb{R}^{QN},\mathcal{V},h_{\alpha})$ defines a sub-Riemannian manifold with a family of bundle metrics for $\alpha\in[0,1]$; see [Che23] for a related discussion and further details.
-2.2. Reparametrization
+2.2. Reparametrization for the $L^{2}$ loss
 Assuming $\operatorname{rank}{D}=QN\leq K$, the adapted gradient flow in output space resulting from (2.26) is simply Euclidean gradient flow,
@@
 Proposition 2.4.
-If $\underline{x}(s)$ is a solution to Euclidean gradient flow (2.47), then
+If $\underline{x}(s)$ is a solution to Euclidean gradient flow (2.46) with respect to the cost (2.47), then
 $\underline{\tilde{x}}(t):=\underline{x}(-N\ln(1-t))=\underline{y}+(1-t)(\underline{x}_{0}-\underline{y}),$ (2.56)
@@
 Solving this matrix valued linear ODE for $R(t)$ using the Duhamel (or variation of constants) formula, and substituting the resulting expression in (2.60) yields (2.58), as claimed. ∎
+2.3. Cross-entropy loss
+For classification of data into $Q$ classes, the output $f_{\underline{\theta}}(x)$ and label $y(x)$ associated to an input $x\in\mathcal{X}_{0}$ are vectors in $\mathbb{R}^{Q}$ representing the probability that $x$ belongs to each class, so that $\ f_{\underline{\theta}}(x)_{j},y_{j}\in[0,1]$, $j=1,\cdots,Q$, and $\sum_{j=1}^{Q}f_{\underline{\theta}}(x)_{j}=\sum_{j=1}^{Q}y_{j}=1.$ The cross-entropy loss is
+$\ell(f(x),y)=-y\cdot\log(f(x))=-\sum_{j=1}^{Q}y_{j}\log(f(x)_{j}),$ (2.64)
+where the $\log$ is taken component-wise on vectors.
+In this case, the task becomes one of comparing the probability distribution of the true labels with that of the predicted labels.
+Prior work has explored the geometric properties of gradient flows on such probability manifolds within the framework of information geometry, see e.g.
+[Ama16, AJVLS17, LM18].
+Whether the ideas presented here can be meaningfully extended to this setting remains an open question, which we leave for future investigation.
+Let us consider instead a model whose output can be any vector in $\mathbb{R}^{Q}$. Let $\sigma$ be the softmax function, so that if $z\in\mathbb{R}^{Q}$, $\sigma(z)\in\mathbb{R}^{Q}$ and
+$\sigma(z)_{j}=\frac{e^{z_{j}}}{\sum_{i}e^{z_{i}}},\quad j=1,\cdots,Q,$ (2.65)
+and define
+$\ell(f(x),y)=-\sum_{j=1}^{Q}y_{j}\log\sigma(f(x))_{j}=-y\cdot\log\sigma(f(x))$ (2.66)
+with $y_{j}\in[0,1]$ for $j=1,\cdots,Q$ and $\sum_{i=1}^{Q}y_{i}=1$. Then Theorem 2.3 applies to
+$\mathcal{C}(\underline{\theta})=\sum_{x\in\mathcal{X}_{0}}\ell(f_{\underline{\theta}}(x),y(x)).$ (2.67)
+Assuming $\operatorname{rank}{D}=QN\leq K$, the adapted gradient flow in output space resulting from (2.26) is simply Euclidean gradient flow,
+$\partial_{t}\underline{x}(t)$ $=-\nabla_{\underline{x}}\mathcal{C}(\underline{x}(t)).$ (2.68)
+From now until the end of this subsection, we hide the dependence on the parameters $\underline{\theta}\in\mathbb{R}^{K}$ and let $f_{t}(x):=f(\underline{\theta}(t),x)$.
+We can compute, for some input $x\in\mathcal{X}_{0}$ and associated label $y=y(x)\in\mathbb{R}^{Q}$,
+$\partial_{f(x)_{k}}\mathcal{C}$ $=\sum_{i\neq k}\left(y_{i}\frac{\sum_{j}e^{f(x)_{j}}}{e^{f(x)_{i}}}\frac{e^{f(x)_{i}}e^{f(x)_{k}}}{\left(\sum_{j}e^{f(x)_{j}}\right)^{2}}\right)-y_{k}\frac{\sum_{j}e^{f(x)_{j}}}{e^{f(x)_{k}}}\frac{\left(\sum_{j}e^{f(x)_{j}}\right)e^{f(x)_{k}}-e^{2f(x)_{k}}}{\left(\sum_{j}e^{f(x)_{j}}\right)^{2}}$
+$=\sum_{i\neq k}\left(y_{i}\frac{e^{f(x)_{k}}}{\sum_{j}e^{f(x)_{j}}}\right)-y_{k}\frac{\sum_{j}e^{f(x)_{j}}-e^{f(x)_{k}}}{\sum_{j}e^{f(x)_{j}}}$
+$=-y_{k}+\frac{e^{f(x)_{k}}}{\sum_{j}e^{f(x)_{j}}}\sum_{i=1}^{Q}y_{i}$ (2.69)
+$=-y_{k}+\frac{e^{f(x)_{k}}}{\sum_{j}e^{f(x)_{j}}},$
+and end up with a system of ordinary differential equations for each input:
+$(f_{t}(x)_{j})^{\prime}=-\partial_{f_{t}(x)_{j}}\mathcal{C}=y_{j}-\frac{e^{f_{t}(x)_{j}}}{\sum_{i}e^{f_{t}(x)_{i}}},\quad j=1,\cdots,Q,$ (2.70)
+or
+$f_{t}(x)^{\prime}=-\nabla_{f_{t}(x)}\mathcal{C}=y-\sigma(f_{t}(x)).$ (2.71)
+We make two key observations: First, note that
+$\sum_{j=1}^{Q}f_{t}(x)_{j}^{\prime}=\underbrace{\sum_{j=1}^{Q}y_{j}}_{=1}-\underbrace{\frac{\sum_{j=1}^{Q}e^{f_{t}(x)_{j}}}{\sum_{i=1}^{Q}e^{f_{t}(x)_{i}}}}_{=1}=0,$ (2.72)
+so
+$\sum_{j=1}^{Q}f_{t}(x)_{j}(s)=\sum_{j=1}^{Q}f_{0}(x)_{j}=:c.$ (2.73)
+Thus, even though we assumed that $f_{t}(x)\in\mathbb{R}^{Q}$ is not necessarily on the probability simplex, so that $c\in\mathbb{R}$ can have any value, a solution to (2.71) is still constrained to an affine hyperplane
+$H_{c}:=\left\{z\in\mathbb{R}^{Q}:z\cdot u_{Q}=c\right\}$ (2.74)
+where $u_{Q}:=(1,\cdots,1)^{T}\in\mathbb{R}^{Q}$.
+Second, we compute the Hessian of the loss (2.66) for a single data point,
+$\left[D^{2}\ell\right]_{jk}$ $=\partial_{f_{t}(x)_{k}}\left(-y_{j}+\frac{e^{f_{t}(x)_{j}}}{\sum_{i}e^{f_{t}(x)_{i}}}\right)$
+$=\delta_{jk}\frac{e^{f_{t}(x)_{j}}}{\sum_{i}e^{f_{t}(x)_{i}}}-\frac{e^{f_{t}(x)_{j}}}{\sum_{i}e^{f_{t}(x)_{i}}}\frac{e^{f_{t}(x)_{k}}}{\sum_{i}e^{f_{t}(x)_{i}}},$ (2.75)
+so
+$D^{2}\ell\,\rvert_{f_{t}(x)}=\operatorname{diag}(\sigma(f_{t}(x)))-\sigma(f_{t}(x))\sigma(f_{t}(x))^{T}.$ (2.76)
+Observe that
+[…]
+Lemma 2.6.
+Let $D^{2}\ell\,\rvert_{z}\in\mathbb{R}^{Q\times Q}$ be the Hessian matrix of the loss (2.66) evaluated at an arbitrary point $z\in\mathbb{R}^{Q}$,
+$D^{2}\ell\,\rvert_{z}=\operatorname{diag}(\sigma(z))-\sigma(z)\sigma(z)^{T}.$ (2.78)
+For all $z\in\mathbb{R}^{Q}$, $\operatorname{rank}D^{2}\ell\,\rvert_{z}=Q-1$ and $D^{2}\ell\,\rvert_{z}$ is positive-semidefinite.
+Proof.
+Fix $z\in\mathbb{R}^{Q}$.
+First, since $\sigma(z)_{j}>0$ for all $j=1,\cdots,Q$, we have $\operatorname{rank}\operatorname{diag}(\sigma(z))=Q$.
+Moreover, $\operatorname{rank}(\sigma(z)\sigma(z)^{T})=1$, so
+$\operatorname{rank}D^{2}\ell\,\rvert_{z}\geq\operatorname{rank}\operatorname{diag}(\sigma(z))-\operatorname{rank}(\sigma(z)\sigma(z)^{T})=Q-1.$ (2.79)
+On the other hand,
+$D^{2}\ell\,\rvert_{z}u_{Q}=0,$ (2.80)
+so $\dim(\ker D^{2}\ell\,\rvert_{z})\geq 1$. Therefore,
+$Q-1\leq\operatorname{rank}D^{2}\ell\,\rvert_{z}=Q-\dim(\ker D^{2}\ell\,\rvert_{z})\leq Q-1.$ (2.81)
+To show $D^{2}\ell\,\rvert_{z}\geq 0$, let $S:=\operatorname{diag}(\sigma(z))$. Then for any $v\in\mathbb{R}^{Q}$,
+$\langle v,D^{2}\ell\,\rvert_{z}v\rangle$ $=\langle v,Sv\rangle-\left\lvert{\langle v,\sigma(z)\rangle}\right\rvert^{2}.$ (2.82)
+Using Cauchy–Schwarz,
+$\left\lvert{\langle v,\sigma(z)\rangle}\right\rvert^{2}$ $=\left\lvert{\langle S^{\frac{1}{2}}v,S^{-\frac{1}{2}}\sigma(z)\rangle}\right\rvert^{2}$ (2.83)
+$\leq\langle v,Sv\rangle\langle\sigma(z),S^{-1}\sigma(z)\rangle.$ (2.84)
+Therefore
+$\langle v,D^{2}\ell\,\rvert_{z}v\rangle\geq\langle v,Sv\rangle\left(1-\langle\sigma(z),S^{-1}\sigma(z)\rangle\right),$ (2.85)
+but
+$S^{-1}\sigma(z)=u_{Q}$ (2.86)
+and
+$\langle\sigma(z),S^{-1}\sigma(z)\rangle=\sigma(z)\cdot u_{Q}=\sum_{j=1}^{Q}\sigma(z)_{j}=1,$ (2.87)
+thus
+$\langle v,D^{2}\ell\,\rvert_{z}v\rangle\geq 0.$ (2.88)
+From this we conclude that $D^{2}\ell\,\rvert_{z}$ is positive-definite when restricted to the submanifolds $H_{c}$, so that for each training data point the cross-entropy loss (2.66) appears convex to a solution of the Euclidean gradient flow (2.71).
+If there exists a critical point, it should correspond to a global minimum.
+Such an equilibrium point must satisfy
+$\sigma(f_{*}(x))=y,$ (2.89)
+and this can only be attained if $y_{j}>0$, for all $j=1,\cdots,Q$111 While it is common to train classifiers using one-hot encoded labels, where $y\in\mathbb{R}^{Q}$ are canonical basis vectors, training with labels that assign positive probabilities to all classes leads to better convergence properties, as we show in the present work.
+This scenario can occur when certain training inputs do not clearly belong to a single class, or can be forced by subjecting the labels to small perturbations.
+We have thus proved most of the following:
+Proposition 2.7.
+If $f_{t}(x)$ is a solution to the gradient flow (2.71) with respect to the cost (2.67), then its image is restricted to an affine hyperplane,
+$\sum_{j=1}^{Q}f_{t}(x)_{j}=c,$ (2.90)
+where $c=\sum_{j=1}^{Q}f_{0}(x)_{j}$.
+Moreover, if the associated label $y\in\mathbb{R}^{Q}$ has positive components, then $f_{t}(x)\to f_{*,c}(x)$ as $t\to\infty$, with
+$f_{*,c}(x)=\log y+\frac{1}{Q}\left(\sum_{j=1}^{Q}(f_{0}(x)_{j}-\log y_{j})\right)u_{Q},$ (2.91)
+where $\log y=(\log y_{1},\cdots,\log y_{Q})^{T}$.
+Note that
+$\nabla_{\underline{x}(t)}\mathcal{C}=(\nabla_{f_{t}(x_{1})}\ell(f_{t}(x_{1}),y(x_{1}))^{T},\cdots,\nabla_{f_{t}(x_{N})}\ell(f_{t}(x_{N}),y(x_{N}))^{T})^{T},$ (2.92)
+and as with the $L^{2}$ loss, the dynamics of training for each data point in the set of training data $\mathcal{X}_{0}$ is independent.
+Further, the Hessian of the total cost (2.67) is block-diagonal,
+$D^{2}\mathcal{C}\,\rvert_{\underline{x}}=\operatorname{diag}(D^{2}\ell\rvert_{f_{t}(x_{1})},\cdots,D^{2}\ell\rvert_{f_{t}(x_{N})})\in\mathbb{R}^{QN\times QN}.$ (2.93)
+It is straightforward to check that $D^{2}\mathcal{C}\,\rvert_{\underline{x}}$ is positive-definite when restricted to a product of hyperplanes $H_{c_{1}}\times\cdots\times H_{c_{N}}\subset\mathbb{R}^{QN}$.
+Therefore, we may extend Proposition 2.7 accordingly.
+Corollary 2.8.
+If $\underline{x}(t)=(f_{t}(x_{1})^{T},\cdots,f_{t}(x_{N})^{T})^{T}\in\mathbb{R}^{QN}$ is a solution to Euclidean gradient flow (2.68) with respect to the cost (2.67), then $\underline{x}(t)$ is constrained to a product of $N$ affine hyperplanes in $\mathbb{R}^{Q}$, defined by
+$\sum_{j=1}^{Q}f_{t}(x_{n})_{j}=c_{n},$ (2.94)
+where $c_{n}=\sum_{j=1}^{Q}f_{0}(x_{n})_{j}$, for $n=1,\cdots,N$.
+Moreover, if the associated label vector $\underline{y}\in\mathbb{R}^{QN}$ has positive components, then there exists a family of global minima $\underline{x}^{*}(C)$ parametrized by $C=(c_{1},\cdots,c_{N})\in\mathbb{R}^{N}$, and $\underline{x}(t)$ converges to a unique equilibrium
+$\underline{x}^{*}(C)$ $=(f_{*,c_{1}}(x_{1})^{T},\cdots,f_{*,c_{N}}(x_{N})^{T})^{T}$ (2.95)
+as $t\to\infty$, for $f_{*,c}$ defined in (2.91).
+Proof.
+All that is left to prove Proposition 2.7 and Corollary 2.8 is to find a formula for $f_{*,c}(x)$. Suppose that $f_{t}(x)$ is such that
+$e^{f_{t}(x)_{j}}=\eta y_{j},\quad j=1,\cdots,Q,$ (2.96)
+so that
+$\sigma(f_{t}(x))_{j}=\frac{e^{f_{t}(x)_{j}}}{\sum_{i}e^{f_{t}(x)_{i}}}=\frac{\eta y_{j}}{\sum_{i}\eta y_{i}}=y_{j}.$ (2.97)
+This satisfies $f_{t}(x)^{\prime}=0$ for (2.71), and is thus an equilibrium point for the gradient flow. Then
+$f_{*,c}(x)_{j}=\log(\eta y_{j})=\log(y_{j})+\log(\eta),\quad j=1,\cdots,Q.$ (2.98)
+If
+$\sum_{j=1}^{Q}f_{0}(x)_{j}=c,$ (2.99)
+then
+$c=\sum_{j=1}^{Q}f_{*,c}(x)_{j}=\sum_{j=1}^{Q}\log y_{j}+Q\log(\eta)$ (2.100)
+and
+$\log(\eta)=\frac{1}{Q}\left(c-\sum_{j}\log y_{j}\right),$ (2.101)
+from which we conclude
+$f_{*,c}(x)_{j}$ $=\log y_{j}+\frac{1}{Q}\left(c-\sum_{i=1}^{Q}\log y_{i}\right)$
+$=\log y_{j}+\frac{1}{Q}\sum_{i=1}^{Q}(f_{0}(x)_{i}-\log y_{i})$ (2.102)
+for each $j=1,\cdots,Q.$ Since $D^{2}\ell\,\rvert_{f_{t}(x)}$ is positive-definite on the hyperplane $H_{c}$, this minimum is unique on $H_{c}$.
+Finally, as
+$\ell(f_{t}(x),y)=-y\cdot\log\sigma(f_{t}(x)),$ (2.103)
+we have
+$\ell(f_{*,c}(x),y)=-y\cdot\log y>0$ (2.104)
+which does not depend on the initial value $f_{0}(x)$. Therefore, $f_{*,c}(x)$ is a degenerate global minimum. ∎
+Remark 2.9.
+For a fixed $x\in\mathcal{X}_{0}$ and two different sets of initial conditions,
+$f_{*,c_{1}}(x)-f_{*,c_{2}}(x)=\frac{1}{Q}(c_{2}-c_{1})u_{Q}$ (2.105)
+or
+$f_{*,c_{1}}(x)=f_{*,c_{2}}(x)+\frac{c_{2}-c_{1}}{Q}u_{Q}.$ (2.106)
+This implies that the equilibrium manifold is a straight line passing through $f_{*,c}(x)$, for any fixed $c\in\mathbb{R}$.
+We conclude that $\mathbb{R}^{Q}$ fibers into invariant manifolds $H_{c}$ for the adapted Euclidean gradient flow, and that the critical manifold $Crit(x)=\left\{z:\nabla_{z}\ell(z,y(x))=0\right\}$ is given by $\left\{f_{*,c}(x)+\lambda u_{Q}:c\in\mathbb{R}\text{ fixed},\lambda\in\mathbb{R}\right\}$.
+Moreover, $Crit(x)$ and $H_{c}$ intersect transversally, for all $c\in\mathbb{R}.$
 3. Applications
@@
 $\Delta\tilde{x}_{j,i_{j}}(t):=\tilde{x}_{j,i_{j}}(t)-\overline{\tilde{x}_{j}}(t)\,,$ (3.10)
-the $\mathcal{L}^{2}$ cost (2.7) decomposes into
+the $L^{2}$ cost (2.47) decomposes into
 $\mathcal{C}[\underline{\tilde{x}}(t)]$ $=\frac{1}{N}\sum_{j=1}^{Q}\sum_{i_{j}}^{N_{j}}|\tilde{x}_{j,i_{j}}(t)-y_{j}|^{2}$
@@
 was supported by NSF grant DMS-2009800 through T.C., and UT Austin’s GS Summer fellowship.
+T.C.
+and P.M.E thank the anonymous referee who suggested we investigate the cross-entropy loss.
 Appendix A Pseudoinverse

### v3

@@
 For the cross-entropy loss, we prove that output space fibers into invariant manifolds over which Euclidean gradient flow converges to a unique global minimum (Proposition 2.8).
+These results underscore the central role of the rank of the Jacobian (or equivalently, the neural tangent kernel) in shaping optimization dynamics.
+In particular, if the gradient flow for a convex loss in output space reaches a fixed point, then at least one of the following must be true: (i) the solution interpolates the data, or (ii) the Jacobian is not injective.
+This becomes evident when considering the constrained Euclidean gradient flow in output space, as is done for the $L^{2}$ loss in section 2.2 and for the cross-entropy loss in section 2.3.
 In section 3, we provide some applications: First, we note that instead of modifying the metric in $\mathbb{R}_{\underline{\theta}}^{K}$, one could prescribe a convenient path in $\mathbb{R}_{\underline{x}}^{QN}$ and then find an associated path in parameter space.
@@
 Finally, we rephrase Theorem 2.3 in terms of the neural tangent kernel.
-We briefly comment on related work in section 4. A short appendix on pseudoinverses for rank deficient matrices is provided as well.
+We briefly comment on related work in section 4. A short appendix on pseudoinverses for rank-deficient matrices is provided as well.
 2. Main results
@@
 It is clear that
-$\left\{\underline{\theta}:\nabla_{\underline{x}(\underline{\theta})}\mathcal{C}=0\right\}\subset\left\{\underline{\theta}:\nabla_{\underline{\theta}}\mathcal{C}=0\right\},$ (2.16)
-however, the two sets are not necessarily identical.
-It is known that if $\operatorname{rank}D[\underline{\theta}]=QN=\min\{K,QN\}$, then $\nabla_{\underline{x}}\mathcal{C}=0$ if, and only if, $\nabla_{\underline{\theta}}\mathcal{C}=0$.
-Moreover, if $\mathcal{C}$ is the squared loss, then $\nabla_{\underline{x}}\mathcal{C}=0$ corresponds to a global minimum with zero loss.
+$\nabla_{\underline{x}(\underline{\theta})}\mathcal{C}=0\implies\nabla_{\underline{\theta}}\mathcal{C}=0,$ (2.16)
+but the converse does not hold in general.
+It is known that if $\operatorname{rank}D[\underline{\theta}]=QN=\min\{K,QN\}$, then $\nabla_{\underline{x}}\mathcal{C}=0\iff\nabla_{\underline{\theta}}\mathcal{C}=0$.
+If $\mathcal{C}$ is the squared loss, then $\nabla_{\underline{x}}\mathcal{C}=0$ corresponds to a global minimum with zero loss.
 See e.g.
-[Coo18, TKB19, Che23, KMTM24, CE25] for more details.
+[TKB19, Coo21, Che23, KMTM24, CE25].
+More generally, consider a cost function for which the only minima in output space are global minima where the model has interpolated the data, $f(\underline{\theta}_{*},x)=y(x)$ for all $x\in\mathcal{X}_{0}$; this is the case for the $L^{2}$ and cross-entropy losses.
+Then for an overparametrized model ($K\geq QN$), we can read the following from (2.15): If the corresponding gradient flow reaches a fixed point, then at least one of the following holds,
+  • the model interpolates the data, or
+  • $D$ is not full-rank.
+In other words, if the interpolating solution is not attainable, gradient flow will only converge when $D$ loses rank.
 Notation.
@@
 Moreover, given a subspace $V\subset\mathbb{R}^{d}$, we denote by $\mathcal{P}_{V}$ the orthogonal projection onto $V$, such that $\mathcal{P}_{V}=\mathcal{P}_{V}^{2}=\mathcal{P}_{V}^{T}$.
+In this work, $V$ will usually be the range (or image) of a linear operator.
 2.1. Adapted gradient flow
@@
 for $\psi$ satisfying
-$DD^{T}\psi=\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C},$ (2.21)
+$DD^{T}\psi=\mathcal{P}_{\operatorname{range}(DD^{T})}\nabla_{\underline{x}}\mathcal{C},$ (2.21)
 results in
-$\partial_{s}\underline{x}(s)=-\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}.$ (2.22)
+$\partial_{s}\underline{x}(s)=-\mathcal{P}_{\operatorname{range}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}.$ (2.22)
 We will now condense these two separate cases into the same. First, note that
@@
 $=DD^{T}(DD^{T}){}^{+}\nabla_{\underline{x}}\mathcal{C}$
-$=\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}.$
+$=\mathcal{P}_{\operatorname{range}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}.$
 Moreover, we can rewrite
@@
 implies
-$\partial_{s}\underline{x}=-\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}=-\mathcal{P}_{\operatorname{ran}(D)}\nabla_{\underline{x}}\mathcal{C}.$ (2.27)
+$\partial_{s}\underline{x}=-\mathcal{P}_{\operatorname{range}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}=-\mathcal{P}_{\operatorname{range}(D)}\nabla_{\underline{x}}\mathcal{C}.$ (2.27)
 We will refer to (2.26) as adapted gradient flow in $\mathbb{R}_{\underline{\theta}}^{K}$ and (2.27) as constrained Euclidean gradient flow in $\mathbb{R}_{\underline{x}}^{QN}$.
+[^1] Note that this is not the same as the Projected Gradient Descent algorithm for optimization on a feasible set.
+In (2.27) the direction of the flow is constrained, not the value of $\underline{x}$.
 It was further shown in [Che23, Theorem 4.2] that the standard gradient flow (2) and the adapted gradient flow (2.26) have the same equilibrium points.
@@
 This holds for any loss function $\ell$ in (2.7), and to any neural network whose activation function has Lipschitz continuous derivatives.
-Non-smooth activation functions, such as ReLU, can be mollified to satisfy this condition.
+Commonly used non-smooth activation functions, such as ReLU, can be smoothed to satisfy this condition, although the resulting gradient flow might depend on the smoothing.
 Proof.
@@
 $V_{\underline{\theta},1}=-(D^{T}D){}^{+}\nabla_{\underline{\theta}}\mathcal{C}.$ (2.33)
-For $\alpha\in[0,1)$, note that $A_{\underline{\theta},\alpha}>0$ for all $\underline{\theta}$. For $\alpha=1$, writing $\mathcal{P}_{\operatorname{ran}(D^{T})}=(D){}^{+}D$, we have
+For $\alpha\in[0,1)$, note that $A_{\underline{\theta},\alpha}>0$ for all $\underline{\theta}$. For $\alpha=1$, writing $\mathcal{P}_{\operatorname{range}(D^{T})}=(D){}^{+}D$, we have
 $\nabla_{\underline{\theta}}\mathcal{C}=(D){}^{+}D\nabla_{\underline{\theta}}\mathcal{C},$ (2.34)
-since (2.15) shows $\nabla_{\underline{\theta}}\mathcal{C}\in\operatorname{ran}(D^{T})$, and also
-$\ker(D^{T}D){}^{+}=\ker(D^{T}D)^{T}=\ker(D^{T}D)=\ker((D){}^{+}D),$ (2.35)
+since (2.15) shows $\nabla_{\underline{\theta}}\mathcal{C}\in\operatorname{range}(D^{T})$, and also
+[…]
 where we used (A.11) and Lemma A.2.
@@
 $=-\left(\alpha D(D^{T}D){}^{+}D^{T}+(1-\alpha)DD^{T}\right)\nabla_{\underline{x}}\mathcal{C}$ (2.38)
-=$\stackrel{{\scriptstyle(*)}}{{=}}-\left(\alpha\mathcal{P}_{\operatorname{ran}(DD^{T})}+(1-\alpha)DD^{T}\right)\nabla_{\underline{x}}\mathcal{C}$
-$=-\left(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)DD^{T}\right)\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C},$
-where ()(*)( ∗ ) follows from (A.7), (A.8) and (A.12). Therefore,
+$\stackrel{{\scriptstyle(*)}}{{=}}-\left(\alpha\mathcal{P}_{\operatorname{range}(DD^{T})}+(1-\alpha)DD^{T}\right)\nabla_{\underline{x}}\mathcal{C}$
+$=-\left(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)DD^{T}\right)\mathcal{P}_{\operatorname{range}(DD^{T})}\nabla_{\underline{x}}\mathcal{C},$
+where ()(*) follows from (A.7), (A.8) and (A.12). Therefore,
 $\partial_{s}\underline{x}(s)=V_{\underline{x}(s),\alpha}$ (2.39)
 for vector fields in $\mathbb{R}^{QN}$ defined along the path $\underline{x}(s)$,
-$V_{\underline{x}(s),\alpha}:=-\left(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)D(s)D(s)^{T}\right)\mathcal{P}_{\operatorname{ran}(D(s)D(s)^{T})}\nabla_{\underline{x}}\mathcal{C}.$ (2.40)
+$V_{\underline{x}(s),\alpha}:=-\left(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)D(s)D(s)^{T}\right)\mathcal{P}_{\operatorname{range}(D(s)D(s)^{T})}\nabla_{\underline{x}}\mathcal{C}.$ (2.40)
 Observe that
@@
 and
-$V_{\underline{x}(s),1}=-\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C},$ (2.42)
+$V_{\underline{x}(s),1}=-\mathcal{P}_{\operatorname{range}(DD^{T})}\nabla_{\underline{x}}\mathcal{C},$ (2.42)
 so that in fact $V_{\underline{x}(s),\alpha},\,\alpha\in[0,1],$ interpolates between the standard dynamics given by (2) at $\alpha=0$, and constrained Euclidean gradient flow (2.27) for $\underline{x}(s)$ at $\alpha=1$.
-Moreover, for $\alpha\in(0,1]$, $(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)DD^{T})$ is positive-definite, and for $\alpha=0$, $DD^{T}$ is injective in the range of $\mathcal{P}_{\operatorname{ran}(DD^{T})}$, so
-$(\alpha{\textbf{1}}+(1-\alpha)DD^{T})\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}=0\text{ if, and only if, }\mathcal{P}_{\operatorname{ran}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}=0.$ (2.43)
+Moreover, for $\alpha\in(0,1]$, $(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)DD^{T})$ is positive-definite, and for $\alpha=0$, $DD^{T}$ is injective in the range of $\mathcal{P}_{\operatorname{range}(DD^{T})}$, so
+$(\alpha{\textbf{1}}+(1-\alpha)DD^{T})\mathcal{P}_{\operatorname{range}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}=0\text{ if, and only if, }\mathcal{P}_{\operatorname{range}(DD^{T})}\nabla_{\underline{x}}\mathcal{C}=0.$ (2.43)
 Thus, the equilibrium points of (2.39) are the same for all $\alpha\in[0,1]$. ∎
@@
 In this situation, $V_{\underline{x}(s),\alpha}$ is the gradient with respect to the bundle metric $h_{\alpha}$ on $\mathcal{V}$ with tensor
-$[h_{\alpha}]=\left(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)DD^{T}\right)^{-1}\Big{|}_{\operatorname{ran}(DD^{T}[\underline{\theta}])}\,.$ (2.45)
+$[h_{\alpha}]=\left(\alpha{\textbf{1}}_{QN\times QN}+(1-\alpha)DD^{T}\right)^{-1}\Big|_{\operatorname{range}(DD^{T}[\underline{\theta}])}\,.$ (2.45)
 The triple $(\mathbb{R}^{QN},\mathcal{V},h_{\alpha})$ defines a sub-Riemannian manifold with a family of bundle metrics for $\alpha\in[0,1]$; see [Che23] for a related discussion and further details.
@@
 and we see that $\underline{x}(s)\to\underline{y}$ as $s\to\infty$.
+Remark 2.4.
+This implies that $\underline{y}$ is in the image of $\underline{x}$, which is not necessarily the case in general.
+Recall from Remark 2.1 that if the interpolating solution is not attainable, then $D$ must lose rank to reach a fixed point.
+Equation (2.49) only holds if $D(t)$ is full rank for all $t\in[0,s]$ and the initialization $\underline{\theta}(0)$.
 Suppose instead that we wanted to impose a linear interpolation
@@
 $\partial_{t}\underline{\tilde{x}}(t)=-\frac{1}{1-t}\mathcal{P}_{t}(\underline{\tilde{x}}(t)-\underline{y})\,,\;\;\;\underline{\tilde{x}}(0)=\underline{x}_{0}\,,\;\;\;t\in[0,1)\,$, (2.57)
-where $\underline{\tilde{x}}(t)=\underline{x}[\underline{\theta}(s(t))]$ and $\mathcal{P}_{t}:=\mathcal{P}_{\operatorname{ran}(DD^{T})}[\underline{\theta}(s(t))]$.
+where $\underline{\tilde{x}}(t)=\underline{x}[\underline{\theta}(s(t))]$ and $\mathcal{P}_{t}:=\mathcal{P}_{\operatorname{range}(DD^{T})}[\underline{\theta}(s(t))]$.
 When $\operatorname{rank}(D)<QN$, the deviation from linear interpolation is given by
-[…]
+$\underline{\tilde{x}}(t)-\big((1-t)\underline{x}_{0}+t\underline{y}\big)=\int_{0}^{t}dt^{\prime}\,\mathcal{U}_{t,t^{\prime}}\,\frac{1-t}{1-t^{\prime}}\,\mathcal{P}^{\perp}_{t^{\prime}}\,(\underline{x}_{0}-\underline{y}),$ (2.58)
 where the linear propagator $\mathcal{U}_{t,t^{\prime}}$ is determined by
@@
 $\partial_{t}R(t)=\frac{1}{1-t}\mathcal{P}_{t}^{\perp}R(t)+\frac{1}{1-t}\mathcal{P}_{t}^{\perp}(\underline{x}_{0}-\underline{y})\,.$ (2.63)
-Solving this matrix valued linear ODE for $R(t)$ using the Duhamel (or variation of constants) formula, and substituting the resulting expression in (2.60) yields (2.58), as claimed. ∎
+Solving this matrix valued linear ODE for $R(t)$ using the Duhamel (or variation of constants) formula (see e.g., [Tes12]), and substituting the resulting expression in (2.60) yields (2.58), as claimed.
+∎
 2.3. Cross-entropy loss
@@
 $\mathcal{C}(\underline{\theta})=\sum_{x\in\mathcal{X}_{0}}\ell(f_{\underline{\theta}}(x),y(x)).$ (2.67)
-Assuming $\operatorname{rank}{D}=QN\leq K$, the adapted gradient flow in output space resulting from (2.26) is simply Euclidean gradient flow,
+Assuming $\operatorname{rank}{D}=QN\leq K$, the gradient flow in output space resulting from (2.26) is simply Euclidean gradient flow,
 $\partial_{t}\underline{x}(t)$ $=-\nabla_{\underline{x}}\mathcal{C}(\underline{x}(t)).$ (2.68)
@@
 $H_{c}:=\left\{z\in\mathbb{R}^{Q}:z\cdot u_{Q}=c\right\}$ (2.74)
-where $u_{Q}:=(1,\cdots,1)^{T}\in\mathbb{R}^{Q}$.
+where $u_{Q}:=(1,\cdots,1)^{T}\in\mathbb{R}^{Q}$. This is in line with the fact that the softmax function is invariant under addition of a scalar multiple of $u_{Q}$.
 Second, we compute the Hessian of the loss (2.66) for a single data point,
@@
 Observe that
-[…]
+$D^{2}\ell\,\rvert_{f_{t}(x)}\,u_{Q}=\sigma(f_{t}(x))-\sigma(f_{t}(x))\underbrace{(\sigma(f_{t}(x))\cdot u_{Q})}_{=\sum_{j}\sigma(f_{t}(x))_{j}\,=\,1}=0.$ (2.77)
 Lemma 2.7.
@@
 $\sigma(f_{*}(x))=y,$ (2.89)
-and this can only be attained if $y_{j}>0$, for all $j=1,\cdots,Q$111 While it is common to train classifiers using one-hot encoded labels, where $y\in\mathbb{R}^{Q}$ are canonical basis vectors, training with labels that assign positive probabilities to all classes leads to better convergence properties, as we show in the present work.
+and this can only be attained if $y_{j}>0$, for all $j=1,\cdots,Q$.
+[^2] While it is common to train classifiers using one-hot encoded labels, where $y\in\mathbb{R}^{Q}$ are canonical basis vectors, training with labels that assign positive probabilities to all classes leads to better convergence properties, as we show in the present work.
 This scenario can occur when certain training inputs do not clearly belong to a single class, or can be forced by subjecting the labels to small perturbations.
@@
 where $\log y=(\log y_{1},\cdots,\log y_{Q})^{T}$.
+Remark 2.9.
+As in the previous section, this corresponds to an interpolating solution. See Remarks 2.1 and 2.4.
 Note that
@@
 In general, from (2.14),
-$\mathcal{P}_{\operatorname{ran}(D^{T})}\partial_{s}\underline{\theta}=(D){}^{+}\partial_{s}\underline{x},$ (3.1)
-and so assuming $\partial_{s}\underline{\theta}=\mathcal{P}_{\operatorname{ran}(D^{T})}\partial_{s}\underline{\theta}$, [^2] Note that this is satisfied by the standard gradient flow $\partial_{s}\underline{\theta}=-\nabla_{\underline{\theta}}\mathcal{C}=-D^{T}\nabla_{\underline{x}}\mathcal{C}.$ we can write
+$\mathcal{P}_{\operatorname{range}(D^{T})}\partial_{s}\underline{\theta}=(D){}^{+}\partial_{s}\underline{x},$ (3.1)
+and so assuming $\partial_{s}\underline{\theta}=\mathcal{P}_{\operatorname{range}(D^{T})}\partial_{s}\underline{\theta}$, [^3] Note that this is satisfied by the standard gradient flow $\partial_{s}\underline{\theta}=-\nabla_{\underline{\theta}}\mathcal{C}=-D^{T}\nabla_{\underline{x}}\mathcal{C}.$ we can write
 $\partial_{s}\underline{\theta}(s)=(D){}^{+}\partial_{s}\underline{x}(s).$ (3.2)
@@
 $\underline{x}(s),$ $\quad s\in[0,T],$
-$\partial_{s}\underline{x}(s)$ $\in\operatorname{ran}(D(s)),$ (3.3)
+$\partial_{s}\underline{x}(s)$ $\in\operatorname{range}(D(s)),$ (3.3)
 $\underline{x}(T)$ $=\underline{y},$
@@
 for some initialization $\underline{\theta}_{0}\in\mathbb{R}^{K}$. Observe that
-$\operatorname{ran}(D)=(\ker(D^{T}))^{\perp}=(\ker(D){}^{+})^{\perp},$ (3.6)
+$\operatorname{range}(D)=(\ker(D^{T}))^{\perp}=(\ker(D){}^{+})^{\perp}$, (3.6)
 and so all equilibrium points of (3.1) satisfy $\partial_{s}\underline{x}=0$.
 As an example, motivated by the results in section 2.2, one could take $\underline{x}(s)$ to be the linear interpolation (2.50) and check if, for a given $\underline{\theta}_{0}$,
-$(\underline{x}_{0}-\underline{y})\in\operatorname{ran}{D(s)},$ (3.7)
+$(\underline{x}_{0}-\underline{y})\in\operatorname{range}{D(s)},$ (3.7)
 for all $s\in[0,T]$.
@@
 Whether or not a path $\underline{x}(s)$ satisfies the constraint in (3.1) (or (3.7) in the example above) could depend on the choice of $\underline{\theta}_{0}$.
-In general, properties of the Jacobian matrix $D[\underline{\theta}(s)]$ that determine the dynamics of $\underline{x}(s)$, such as rank and the subspace $\operatorname{ran}(DD^{T})$, might depend on the initialization even within a class $\left\{\underline{\theta}_{0}:\underline{x}(\underline{\theta}_{0})=\underline{x}_{0}\right\}$.
+In general, properties of the Jacobian matrix $D[\underline{\theta}(s)]$ that determine the dynamics of $\underline{x}(s)$, such as rank and the subspace $\operatorname{range}(DD^{T})$, might depend on the initialization even within a class $\left\{\underline{\theta}_{0}:\underline{x}(\underline{\theta}_{0})=\underline{x}_{0}\right\}$.
 3.2. Final layer collapse
@@
 $\sum_{i=1}^{N}\left\lvert{u_{i}}\right\rvert^{2}>0\text{ implies }\sum_{i,j=1}^{N}u_{i}^{T}\Theta(x_{i},x_{j},\underline{\theta})u_{j}>0\,.$ (3.16)
-From (3.15) we see that his holds if, and only if, $DD^{T}$ is a positive-definite matrix.
+From (3.15) we see that this holds if, and only if, $DD^{T}$ is a positive-definite matrix.
 Thus, Theorem 2.3 suggests that, by changing the metric in parameter space, the NTK can be modified into a simpler kernel $K$ which preserves the equilibrium points of the flow.
@@
 4. Some related work
+Preconditioners
+The preconditioner $(D^{T}D)^{+}$ introduced here closely resembles the inverse of the generalized Gauss–Newton matrix (GN), which approximates the Hessian of the loss with respect to the parameters:
+$\left(D^{T}D\right)_{ij}$ $=\,\,\,\,(\partial_{i}\underline{x}^{T})(\partial_{j}\underline{x}),$
+$\frac{\partial^{2}\mathcal{C}}{\partial\theta_{i}\partial\theta_{j}}\,\,\,$ $=\underbrace{(\partial_{i}\underline{x}^{T})H_{\underline{x}}(\partial_{j}\underline{x})}_{(GN)_{ij}}+\partial_{i}\partial_{j}\underline{x}^{T}\nabla_{\underline{x}}\mathcal{C}.$ (4.1)
+The distinction lies in omitting $H_{\underline{x}}$, the Hessian with respect to the outputs, which is the identity for the $L^{2}$ loss.
+The proof of Theorem 2.3 remains valid for $GN^{-1}$ as long as $H_{\underline{x}}$ is positive-semidefinite in the region of interest (which holds for the $L^{2}$ and cross-entropy losses), but fails in general.
+Recent work has examined the properties of the GN matrix [ZSL24], as well as potential advantages [AVKM25] of the full Gauss–Newton matrix compared to Adam [KB17].
+For a broader discussion on curvature-aware optimization, see e.g.
+[Mar20] and references therein.
 Linear paths
@@
 [COB19, WGL+20].
-Certain works study the rank of $D$ more explicitly: In [GLMW22], the authors study the functional dimension of ReLU neural networks, defined in terms of the rank of the Jacobian.
+Certain works study the rank of $D$ more explicitly: In [GLMW25], the authors study the functional dimension of ReLU neural networks, defined in terms of the rank of the Jacobian.
 In [KMTM24], the authors show that for mildly overparametrized ReLU neural networks with scalar output ($Q=1$) most activation patterns correspond to regions in parameter space where the Jacobian is full rank.
@@
 T.C.
-and P.M.E thank the anonymous referee who suggested we investigate the cross-entropy loss.
+and P.M.E.
+thank the anonymous referee who suggested we investigate the cross-entropy loss.
 Appendix A Pseudoinverse
@@
 Then
-$A(A){}^{+}$ $=\mathcal{P}_{\operatorname{ran}(A)},$ (A.8)
-$(A){}^{+}A$ $=\mathcal{P}_{\operatorname{ran}(A^{T})}.$ (A.9)
+$A(A){}^{+}$ $=\mathcal{P}_{\operatorname{range}(A)},$ (A.8)
+$(A){}^{+}A$ $=\mathcal{P}_{\operatorname{range}(A^{T})}.$ (A.9)
 When considering the subspaces associated to $A$, it is useful to know the relations
-$\operatorname{ran}(A){}^{+}$ $=\operatorname{ran}A^{T},$ (A.10)
+$\operatorname{range}(A){}^{+}$ $=\operatorname{range}A^{T},$ (A.10)
 $\ker(A){}^{+}$ $=\ker A^{T}.$ (A.11)
@@
 Lemma A.2.
-$\ker(A^{T}A)=\ker((A){}^{+}A)$.
+[…]
 Proof.
@@
 and so
-$\ker(A^{T}A)\subset\ker((A){}^{+}A).$ (A.16)
+[…]